OpenStudy (nali):

find all solutions of the equation in the interval [0,2pi) sin(x + pi/2 ) - sin (x- pi/2) =3^1/3

OpenStudy (nali):

thanks i got the answer

OpenStudy (anonymous):

\[ \sin (x)-\sin (y)=2 \sin \left(\frac{x-y}{2}\right) \cos \left(\frac{x+y}{2}\right) \] Did you use this formula?

OpenStudy (nali):

no i used this equation sin (u+v) = sin u cos v + cos u sin v

OpenStudy (anonymous):

That will do. The formula I mentioned gives you \[ 2 \cos(x) = \sqrt 3\\ x=\pm \frac {\pi}6+ 2 k \pi \] Did you get this answer?

OpenStudy (nali):

yes thanku

OpenStudy (anonymous):

yw